OpenStudy (anonymous):

Factor the trinomial: 3x2 + 11x + 6

OpenStudy (anonymous):

Is there a factor that is common in each term?

OpenStudy (anonymous):

A. (3x + 2)(x + 3) B. (3x + 3)(x + 2) C. (3x + 1)(x + 6) D. (3x + 6)(x + 1) these are the options

OpenStudy (anonymous):

Okay, so basically, we need to find two binomials then when multiplied, they give us the original equation.

OpenStudy (anonymous):

So do you know how to foil?

OpenStudy (anonymous):

@Haiz, Im trying to teach how to do it, not give the answers!

OpenStudy (anonymous):

yes but im not so good at it

OpenStudy (anonymous):

Okay, so we need to find two binomials. So we start by setting up the following: \(\ \Huge (3x + || )(x + ||) \). We know that the first terms have to equal \(\ 3x^2 \). 3x and x multiplied together equal 3x^2. Next, we know that the outer term has to equal 6. what numbers could we plug in for the space (indicated by the ||) that would equal 6 when multiplied?

OpenStudy (anonymous):

the standard equation is ax2+bx+c here we have 3x2 + 11x + 6 now compare the two equations a = 3, b =11, c =6 now find the product a*c= 3*6=18 now factor 18 such that by addition or subtraction of the factors u get b = 11. which is possible as 18=9*2 (9+2=11) so we can write the equation as 3x2 + 11x + 6 3x2 + 9x+2x + 6 bracket the terms to make easier (3x2 + 9x)+(2x + 6) = 3x (x+3)+2(x+3) =(3x+2)(x+3) hope u got it..